A box contains

5

plain pencils and

5

pens. A second box contains

7

color pencils and

3

crayons. One item from each box is chosen at random. What is the probability that a pen from the first box and a crayon from the second box are selected?

Verified Answer :

5:5 (first box, pencils to pens)

7:3 (second box, coloured pencils to crayons)

The probability of picking a pen (1st box): 5/10

The probability of picking a crayon (2nd box): 3/10

Probability of picking both: 5/10*3/10 = 15/100


the pen has a 50% probability.The crayon has 4.28 probability.